Twenty years ago Cody invested $8,000 in a savings bond. If the bond matures today for $71,970 what was its annual rate of change? Note: Please make sure your final answer(s) are in percentage form and are accurate to 2 decimal places. For example 34.56%
Answer #1

For calculating annual rate of change, we need to find out 20 years rate of change.

That is = 71,970/8,000*100 =899.63%

Annual rate of change = 899.63/20 =44.98%